BeautifulAngel:

A box contains 20 lightbulbs, of which 5 are defective. If 4 lightbulbs are picked from the box randomly, the probability that at most 2 of them are defective is ?

BeautifulAngel:

1 attachment
BeautifulAngel:

@hero

satellite73:

at most two means 0, 1, or 2 you have to compute each of these separately and add them up

satellite73:

the probability that none are defective would be \[\frac{15}{20}\times \frac{14}{19}\times \frac{13}{18}\times \frac{12}{17}\] there may be a more elegant way to do this problem but i can't think of it or maybe do \[\frac{\binom{15}{4}}{\binom{20}{4}}\]

satellite73:

you seen any of this before?

BeautifulAngel:

No, I don't understand math at all. That's why I'm repeating the class.

satellite73:

this question is quite involved, it will take a lot of doing are you up for it?

BeautifulAngel:

Sure. I literally have no choice if i want to be successful, right?

satellite73:

ok so first, have you ever seen this notation before ?

satellite73:

\[\frac{\binom{15}{4}}{\binom{20}{4}}\]

BeautifulAngel:

No, never.

satellite73:

how about this \[_{15}C_4\]

satellite73:

it is read "fifteen choose four" are you familiar with that ?

BeautifulAngel:

I've seen it before but I've never used it. This is all new to me.

satellite73:

it is the number of ways to choose 4 out of a set of 15 if you have 20 light bulbs, 5 of which are defective, then there are 15 that are not defective that is clear right?

satellite73:

this is where you say "yes, that is obvious"

BeautifulAngel:

yes, sorry. I wasn't at my laptop

satellite73:

ok so the question is, if there are 15 that are not defective, how many ways can you choose 4 of them the answer is 15 choose 4 computed via \[_{15}C_4=\frac{15\times 14\times 13\times 12}{4\times 3\times 2}\] it is a whole number, so cancel first, multiply last or else use wolfram

satellite73:

http://www.wolframalpha.com/input/?i=(15+choose+4)

BeautifulAngel:

Okay. You've completely lost me. I'm sorry, I've been up for days and working at this for hours. I have to have it done by the 30th this month so late nights, early morning, or just plain no sleep at all. I just never got algebra anyway. Anyway, you've completely lost me there.

satellite73:

then i suggest you give this a rest, because this is only one tiny part of the question, which, as i said, is fairly involved you have to compute \({15}C_4\) but also \({20}C_4\) and that is just for the probability you get 0 defective ones, so that is only a start

satellite73:

by the way, this is not algebra, it is probability, which involves mostly arithmetic if you understand how to do the problem

BeautifulAngel:

I'd love some sleep, but I'm tired of my family griping about my "failing classes and grades". But thank you for helping with what you could. I'll tag you in the morning for some more help if that's okay.

satellite73:

sure and good luck just for my own information, how were you going to try to solve this question?

BeautifulAngel:

I have no idea, but all the answers that I guessed on were wrong and it doesn't give me any other choices. the answers they gave me to choose from were: 31/969 70/323 91/323 938/969

satellite73:

then lets just do it, it can't take more than ten minutes

BeautifulAngel:

Okay, that works. Thank you

satellite73:

the total number of ways to pick 4 out of 20 is called 20 choose 4 the total number of ways to pick 4 out of 15 is called 15 chose 4 to the probability you get no defective ones is 15 choose 4 over 20 choose 4 forget about how to compute this, use technology http://www.wolframalpha.com/input/?i=(15+choose+4)%2F(20+choose+4)

satellite73:

copy and paste that, we need that number now we compute the number of ways to get 1 defective one the denominator will still be 20 choose 4 the numerator will be 15 choose 3 times 5 chose 1 (3 not defective, one defective)

satellite73:

here is that number http://www.wolframalpha.com/input/?i=(15+choose+3)*(5+choose+1)%2F(20+choose+4)

satellite73:

now the probabilty you get two defecttive 15 choose 2 times 5 chose 2 over 20 choose 4

satellite73:

here is that number http://www.wolframalpha.com/input/?i=(15+choose+3)*(5+choose+1)%2F(20+choose+4)

BeautifulAngel:

Hey, this is her teacher. Even I don't know what I'm doing, could you please just give us the answer so this can be over with. (Also, the mother.)

satellite73:

now we have to add them up so lets let wolfram do it http://www.wolframalpha.com/input/?i=((15+choose+4)%2B(15+choose+3)(5+choose+1)+%2B(15+choose+2)(5+choose+2))%2F(20+choose+4)

satellite73:

copy and paste, the answer is there

satellite73:

teacher?

BeautifulAngel:

Yes. I'm "BeautifulAngel's" teacher and mother.

BeautifulAngel:

What am I copying and pasting?

satellite73:

http://www.wolframalpha.com/input/?i=((15+choose+4)%2B(15+choose+3)(5+choose+1)+%2B(15+choose+2)(5+choose+2))%2F(20+choose+4) home school?

Ultrilliam:

Note to self, fix the url parser so it does not break with ('s.

satellite73:

i think that can't be helped

satellite73:

by which i mean i think it did the same thing in OS

BeautifulAngel:

No, she goes to a public school. But I teach senior honors and cambridge math. She goes to school in Tennessee.

satellite73:

wolfram says \[\frac{938}{969}\] which seems high to me, but whatever

Ultrilliam:

I can fix it though lol I just didn't account for it while writing the function that verifies a url

satellite73:

don't know cambridge math heard of cambridge though...